OpenStudy (anonymous):

simplify (10r+9)-(12r-5)

OpenStudy (anonymous):

see here u have (10r + 9 ) - (12r - 5) = 10r + 9 - 12r + 5 = 10r - 12r + 9 - 5 = -2r + 4 take 2 as common in -2r + 4 expression u would get: 2(-1r + 2) or -2(r + 2)

OpenStudy (anonymous):

hope it helped eva 147

OpenStudy (anonymous):

it actually did thanks alot

OpenStudy (aroub):

10r+9-12r+5 = -2r+14

OpenStudy (anonymous):

You can just think of it like this.. From 10 r you are subtracting 12 r, and from 9 ones you are subtracting -5 ones. Just do the operation for each type separately, and the result will be the sum of the results of each.

OpenStudy (anonymous):

i did mistake here so sorry it would be -2(r - 7) or 2(-r + 7)

OpenStudy (anonymous):

-2r+4